Hong is standing on a hill and throws a frisbee to her friend, who stands on flat ground at the bottom of the hill. The toss is short, however, and the frisbee hits the ground before her friend can get to it. The height of the frisbee, in meters, above the flat ground at the bottom of the hill, y, is dependent on the number of seconds after it is thrown, x, and can be modeled with the function y=-(x-2)^2+36
How many seconds after being thrown will the frisbee reach its maximum height? .. seconds

What is the maximum height that the frisbee will reach?....meters

How many seconds after being thrown will the frisbee hit the ground? .seconds..

What is the height of the frisbee at the time it is thrown? ...meters

Answer by stanbon(75887)

The height of the frisbee, in meters, above the flat ground at the bottom of the hill, y, is dependent on the number of seconds after it is thrown, x, and can be modeled with the function
y=-(x-2)^2+36
How many seconds after being thrown will the frisbee reach its maximum height?
Vertex:: (2,36)
Ans: 2 seconds
----------------------
What is the maximum height that the frisbee will reach?....meters
Ans: f(2) = 36
-------------------------
How many seconds after being thrown will the frisbee hit the ground?
Solve -(x-2)^2+36 = 0
(x-2)^2 = 36
x-2 = 6
x = 8 seconds
-----------------------
What is the height of the frisbee at the time it is thrown? ...meters
f(0) = -(0-2)^2+36
f(0) = -4+36
Ans: 32 meters
----------------------
Cheers,
Stan H.
-----------------